Clairaut's theorem holds that Uxy=Uyx,show that the following equations obey clairaut's theorem.

1.u=in(x+2y).

2.u=e^xy siny.


1
Expert's answer
2021-06-28T16:29:43-0400

1)"U = ln(x+2y)"

"U_x = \frac{1}{x+2y}\\
U_{xy} = \frac{-2}{(x+2y)^2}"

Also,

"U_y = \frac{2}{x+2y}\\
U_{yx} = \frac{-2}{(x+2y)^2}"

Clairaut's theorem as being justified as desired

2) "U = e^{xy} \sin y"

"U_x = ye^{xy}\sin y \\ U_{xy} = e^{xy}\sin y + + xy e^{xy}\sin y + ye^{xy}\cos y"

"U_y = xe^{xy}\sin y + e^{xy}\cos y\\ U_{yx} = e^{xy}\sin y + + xy e^{xy}\sin y + ye^{xy}\cos y"

Clairaut's theorem as being justified as desired.